OpenStudy (anonymous):

The equation of a line that goes through the points (5,9) and(6,3) can be written in the form y=mx+b. Where is m? Where is b?

OpenStudy (jlastino):

You should first make that equation into the slope intercept form y=mx+b -5x+3y+6=0 3y=5x-6 Divide everything by 3 What would be the new equation?

OpenStudy (jlastino):

Dude your slope is 5/3 Just express that in decimal

OpenStudy (jlastino):

Have you tried inputting 1.67? because I think it won't accept 1.666666667....

OpenStudy (jlastino):

Ok for the new question use the slope-point formula y-y1= m(x-x1) y-3=4(x-7) y-3 =4x-28 y=4x-25 m=4 b=-25

OpenStudy (jlastino):

use the slope-point formula (again :)) ) m=-2 P(6,-4) y-y1= m(x-x1) y+4=-2(x-6) y+4 =-2x+12 y=-2x+8 m=-2 b=8

OpenStudy (jlastino):

You know it would be better if you post your questions individually instead of editing so everyone can see it easier

OpenStudy (campbell_st):

find the gradient( slope). Its rise/run \[m = (y _{2} - y _{1})/(x _{2} - x _{1})\] in your question use (6,3) as the 2 point and (5, 9) as the 1 point then \[m = (3 - 9)/(6 - 5)\] the gradient(slope) of the line is -6 next use the point gradient formula to find the equation of the line \[y - y _{1} = m( x - x _{1})\] using m = -6 and the point ( (5, 9) y - 9 = -6(x - 5) expand and simplify y = -6x + 39